### v1

@@ Abstract
This paper introduces a novel, entity-aware metric, termed as Radiological Report (Text) Evaluation (RaTEScore), to assess the quality of medical reports generated by AI models. RaTEScore emphasizes crucial medical entities such as diagnostic outcomes and anatomical details, and is robust against complex medical synonyms and sensitive to negation expressions. Technically, we developed a comprehensive medical NER dataset, RaTE-NER, and trained an NER model specifically for this purpose. This model enables the decomposition of complex radiological reports into constituent medical entities. The metric itself is derived by comparing the similarity of entity embeddings, obtained from a language model, based on their types and relevance to clinical significance. Our evaluations demonstrate that RaTEScore aligns more closely with human preference than existing metrics, validated both on established public benchmarks and our newly proposed RaTE-Eval benchmark.
@@ 1 Introduction
With the general advancement in nature language processing OpenAI (2023); Anil et al. (2023); Qiu et al. (2024); Wu et al. (2024) and computer vision Li et al. (2023); Alayrac et al. (2022); OpenAI ; Zhang et al. (2023), the pursuit of generalist medical artificial intelligence has grown increasingly promising and appealing Moor et al. (2023); Wu et al. (2023); Tu et al. (2024). Yet, the complexity and specialized nature of clinical free-form texts, such as radiology reports and discharge summaries, present substantial challenges in evaluating medical foundation models.
@@
In the existing literature, four main types of metrics have been adopted to assess the similarity between free-form texts in medical scenarios, as shown in Figure 1. These include: (i) Word Overlap Metrics, such as BLEU Papineni et al. (2002) and ROUGE Lin (2004). While intuitive, these metrics fail to capture negation or synonyms in sentences, thus neglecting the semantic factuality; (ii) Embedding Similarity Metrics, like BERTScore Zhang et al. (2019), provide better semantic awareness but fail to emphasize key medical terms, leading to overlooked errors in critical conclusions; (iii) Metrics based on Named Entity Recognition (NER), such as RadGraph F1 Yu et al. (2023a) and MEDCON Yim et al. (2023). Although tailored for the medical domain, they struggle with synonym unification and are typically restricted to analyzing Chest X-ray reports; (iv) Metrics relying on large language models (LLMs), such as those proposed by Wei et al.Wei et al. (2024) and Liu et al.Liu et al. (2023). While these metrics are better aligned with human preferences, they suffer from potential subjective biases and are prohibitively costly for large-scale evaluation.
@@
Second, we have developed a comprehensive medical named-entity recognition (NER) dataset, RaTE-NER, which encompasses data from 9 modalities and 22 anatomical regions, derived from MIMIC-IV and Radiopaedia. In addition, we introduce RaTE-Eval, a novel benchmark designed to assess metrics across various clinical texts. This benchmark is structured around three sub-tasks: Sentence-level Human Rating, Paragraph-level Human Rating, and Comparison of Synthetic Reports, each targeting specific evaluation challenges. Both the RaTE-NER dataset and the RaTE-Eval benchmark will be made publicly available, aiming to foster the development of more effective evaluation metrics within the field of medical informatics.
Third, our extensive experiments demonstrate the superiority of our proposed RaTEScore. We initially tested our metric against other metrics on the public dataset ReXVal Yu et al. (2023a), it showing superior alignment with human preference. Given ReXVal’s limitation to chest X-ray reports, further testing was conducted on the diverse sub-tasks of RaTE-Eval, where RaTEScore consistently outperformed other metrics. We conducted ablation studies to validate the effectiveness of individual components of our pipeline.
@@ 2 Methods
In this section, we start by properly formulating the problem, and introduce the pipeline of our metric (Sec. 2.1). Then, we detail each of the module for our metric computation, for example, medical named entity recognition (Sec. 2.2), synonym disambiguation encoding (Sec. 2.3), and the final scoring procedure (Sec. 2.4). Lastly, we present the details for training and evaluation at each stage.
@@ 2.1 General Pipeline
Given two radiological reports, one is the ground truth for reference, denoting as $x$, and the other candidate for evaluation as $\hat{x}$. We aim to define a new similarity metric $S(x,\hat{x})$, that enables to compare two radiological reports at the entity level, thus better reflecting their clinical consistency.
@@
$\mathbb{F}=\Phi_{\text{ENC}}(\Phi_{\text{NER}}(x)),$ (1)
@@
$S(x,\hat{x})=\Phi_{\text{SCO}}(\mathbb{F},\mathbb{\hat{F}})$ (2)
@@ 2.2 Medical Named Entity Recognition
Overall, we categorize the entity types into five distinct groups within radiological contexts: {Anatomy, Abnormality, Disease, Non-Abnormality, Non-Disease}. Specifically, ‘Abnormality’ refers to notable radiological features such as masses, effusion, and edema. Conversely, ‘Non-Abnormality’ denotes cases where such abnormalities are negated in the context, as illustrated by the classification of ‘pleural effusion’ in the statement ‘No evidence of pleural effusion’.
@@
RaTE-NER Dataset. To support the development of our medical entity recognition module, we have constructed the RaTE-NER dataset, a large-scale, radiological named entity recognition (NER) dataset, including 13,235 manually annotated sentences from 1,816 reports within the MIMIC-IV database, that spans 9 imaging modalities and 23 anatomical regions, ensuring comprehensive coverage. Given that reports in MIMIC-IV are more likely to cover common diseases, and may not well represent rarer conditions, we further enriched the dataset with 33,605 sentences from the 17,432 reports available on Radiopaedia Rad , by leveraging GPT-4 and other medical knowledge libraries to capture intricacies and nuances of less common diseases and abnormalities. More details can be found in the Appendix A.2. We manually labeled 3,529 sentences to create a test set, as shown in Table 2 and Table 1, the RaTE-NER dataset offers a level of granularity not seen in previous datasets, with comprehensive entity annotations within sentences, that enables to train models for medical entity recognition within our analytical pipeline.
@@ 2.4 Scoring Procedure
$S(x,\hat{x})=\frac{\sum_{j}{W(t_{i^{*}},t_{j}})\cdot\text{sim}(e_{i^{*}},\hat{e}_{j})}{\sum_{j}{W(t_{i^{*}},t_{j})}},$
@@
$\text{sim}(e_{i},\hat{e}_{j})=\left\{\begin{aligned} {p}\cos(f_{i},\hat{f}_{j}),\quad if\quad t_{i}\neq t_{j}\\ \cos(f_{i},\hat{f}_{j}),\quad if\quad t_{i}=t_{j}\end{aligned}\right.,$
@@
Finally, due to the order of performing max indexing and mean pooling, the finial similarity metric $S(x,\hat{x})$ does not comply with the commutative law. $S(x,\hat{x})$ and $S(\hat{x},x)$ can be analogous to precision and recall respectively. Thus, to take care of both, our final RaTEScore is defined following the classical F1-score format, as:
$\texttt{RaTEScore}=2\times\frac{{S(x,\hat{x})}\times{S(\hat{x},x)}}{{S(x,\hat{x})}+{S(\hat{x},x)}}$ (3)
@@ 3 RaTE-Eval Benchmark
Sentences-level Human Rating. Existing studies have predominantly utilized the ReXVal dataset Yu et al. (2023b), which requires radiologist annotators to identify and count errors in various potential categories. The metric’s quality is assessed by the Kendall correlation coefficient between the total number of errors and the result from automatic metric. The possible error categories are as follows:
@@ 5. Mention the comparison that is absent from the reference impression;
Building on this framework, we introduce two improvements to enhance the robustness and applicability of our benchmark: (1) normalization of error counts: recognizing that a simple count of errors may not fairly reflect the informational content in sentences, we have adapted the scoring to annotate the number of potential errors. This approach normalizes the counts, ensuring a more balanced assessment across varying report complexities. (2) diversification of medical texts: unlike existing benchmarks that are limited to Chest x-rays from the MIMIC-CXR dataset Johnson et al. (2019), our dataset includes 2215 reports spanning 9 imaging modalities and 22 anatomies from the MIMIC-IV dataset Johnson et al. (2020), the involving imaging modalities and anatomies are listed in Appendix A.3.
@@
Paragraph-level Human Rating. Given that medical imaging interpretation commonly involves the evaluation of lengthy texts rather than isolated sentences, we also incorporate paragraph-level assessments into our analysis, from the MIMIC-IV reports. However, as it is challenging for humans to completely count all errors in long paragraphs accurately, we established a 5-point scoring system for our evaluations, following the RadPEER Goldberg-Stein et al. (2017)-an internationally recognized standard for radiologic peer review. The scores range from 5, denoting a perfectly accurate report, to 0, that indicates the report without any correct observations. Detailed scoring criteria are provided in Appendix A.4, guiding radiologists on how to assign scores at different levels.
@@ 4.4.2 Entity Encoding Module Discussion
In our entity encoding evaluation, we compare two off-the-shelf entity encoding models on the sentence-level correlation task of RaTE-Eval. The first model, BioLORD-2023-C, is trained on medical entity-definition pairs, while the second, MedCPT-Query-Encoder, is trained on PubMed user click search logs. The models achieved Pearson correlation coefficients of 0.54 and 0.52, respectively. BioLORD outperforms MedCPT with 0.02 in Pearson Consistency. Based on these results, we selected BioLORD-2023-C as the base model for our Entity Encoding Module.
@@ 5.2 Radiological Text Evaluation Metric
With the advancement of medical image analysis, researchers have recognized the importance of evaluating the quality of radiology text generation. Metrics such as CheXbert F1 Smit et al. (2020) and RadGraph F1  Yu et al. (2023a) are based on medical entity extraction models. However, CheXbert can only annotate 14 chest abnormalities, and RadGraph F1 Jain et al. (2021) is only trained on chest X-ray modality. MEDCON Yim et al. (2023) expands the extraction range by QuickUMLS package Soldaini and Goharian (2016), which relies on a string match algorithm that is not flexible. RadCliQ Yu et al. (2023a) performs ensembling with BLEU, BERTScore, CheXbert vector similarity, and RadGraph F1 for a comprehensive yet less interpretable evaluation. These metrics calculate the overlap between reference and candidate sentences while overlooking the issue of synonymy. Recently, metrics using Large Language Models (LLMs) such as GPT-4, such as G-Eval Liu et al. (2023), LLM-as-a-Judge Zheng et al. (2024), and LLM-RadJudge Wang et al. (2024) have emerged, closely mimic human evaluation levels. However, these methods are unexplainable and may have potential subjective bias. Besides, their high computational cost also limits them for statistic robust large-scale evaluation.
@@ 5.3 Medical Named-Entity Recognition
The MedNER task targets extracting medical-related entities from given contexts. Great efforts have been made in this domain Jin et al. (2023); Monajatipoor et al. (2024); Keloth et al. (2024); Li and Zhang (2023); Chen et al. (2023). Inspired by the success of this work, we believe MedNER models are strong enough to simplify complex clinical texts, thus reducing the difficulty of automatically comparing two clinical texts. The most related work to ours is RadGraph Jain et al. (2021) which trained an NER model for Chest X-ray reports while we are targeting the general clinical report regardless of their type.
@@ 6 Conclusion
Although our proposed metric, RaTEScore, has performed well across various datasets, there are still some limitations. First, in the synonym disambiguation module, we evaluated the performance of several existing models and directly ultilized them without fine-tuning specifically for the evaluation scenario, which could be enhanced in the future. Furthermore, while we expanded from single-modality radiological report evaluation to multimodal whole-body imaging, we still only considered the issues within the radiological report scenario and did not extend to other medical contexts beyond radiology, nor to the evaluation of other medical tasks, like medical QA, summarisation task. These areas require ongoing research and exploration.
@@ A.2 Automatic Annotation Approach
Specifically, we manually annotate several reports at the required granularity and adopt few-shot prompts with GPT-4 to initially establish an NER dataset. Following this, we build a robust medical entity library, integrating UMLS Bodenreider (2004), Snomed CT Donnelly et al. (2006), ICD-10 ICD , and other knowledge bases, then, compare all extracted entities using the MedCPT Jin et al. (2023) model for similarity. During the comparison process, entities with cosine similarity lower than 0.83 were filtered out. Most entities below this threshold did not meet our requirements. Subsequently, we removed sentences with an entity annotation density lower than 0.7 at the sentence level. Finally, we use medspaCy Eyre et al. (2021) and also key negative words detection in reports, such as “no”, “without”, “unremarkable”, “intact”, to determine the positive or negative polarity of each word in the sentence.

### v2

@@ Abstract
This paper introduces a novel, entity-aware metric, termed as Radiological Report (Text) Evaluation (RaTEScore), to assess the quality of medical reports generated by AI models. RaTEScore emphasizes crucial medical entities, such as diagnostic outcomes and anatomical details. Moreover, it is robust against medical synonyms and sensitive to negation expressions. Technically, we developed a comprehensive medical NER dataset, RaTE-NER, and trained an NER model specifically for this purpose. This model enables the decomposition of complex radiological reports into constituent medical entities. The metric itself is derived by comparing the similarity of entity embeddings, obtained from a language model, based on their types and relevance to clinical significance. Our evaluations demonstrate that RaTEScore aligns more closely with human preference than existing metrics, validated both on established public benchmarks and our newly proposed RaTE-Eval benchmark.
@@ 1 Introduction
With the general advancement in natural language processing OpenAI (2023); Anil et al. (2023); Qiu et al. (2024); Wu et al. (2024) and computer vision Li et al. (2023); Alayrac et al. (2022); OpenAI ; Zhang et al. (2023); Wu et al. (2023a); Zhou et al. (2024), the pursuit of generalist medical artificial intelligence has grown increasingly promising and appealing Moor et al. (2023); Wu et al. (2023b); Tu et al. (2024); Zheng et al. (2023); Zhao et al. (2023), leading to the development of free-text generative foundation models capable of understanding and interpreting radiology imaging studies. Yet, the complexity and specialized nature of clinical free-form texts, such as radiology reports and discharge summaries, present substantial challenges in evaluating model performance. There is an urgent need for a robust and lightweight free-text evaluation metric to better monitor the development of medical generative foundation models and drive advancements in the field of generalist medical artificial intelligence.
@@
In the existing literature, four main types of metrics have been adopted to assess the similarity between free-form texts in medical scenarios, as shown in Figure 1. These include: (i) Word Overlap Metrics, such as BLEU Papineni et al. (2002) and ROUGE Lin (2004). While intuitive, these metrics fail to capture negation or synonyms in sentences, thus neglecting the semantic factuality; (ii) Embedding Similarity Metrics, like BERTScore Zhang et al. (2020), provide better semantic awareness but fail to emphasize key medical terms, leading to overlooking errors in critical conclusions; (iii) Metrics based on Named Entity Recognition (NER), such as RadGraph F1 Yu et al. (2023a) and MEDCON Yim et al. (2023). Although tailored for the medical domain, they struggle with synonym unification and are typically restricted to analyzing Chest X-ray reports; (iv) Metrics relying on large language models (LLMs) Wei et al. (2024); Liu et al. (2023). While these metrics are better aligned with human preferences, they suffer from potential subjective biases and are prohibitively costly for large-scale evaluation.
@@
Second, we have developed a comprehensive medical named-entity recognition (NER) dataset, RaTE-NER, which encompasses data from 9 modalities and 22 anatomical regions, derived from MIMIC-IV and Radiopaedia. In addition, we introduce RaTE-Eval, a novel benchmark designed to assess metrics across various clinical texts. This benchmark is structured around three sub-tasks: Sentence-level Human Rating, Paragraph-level Human Rating, and Comparison of Synthetic Reports, each targeting specific evaluation challenges. Both the RaTE-NER dataset and the RaTE-Eval benchmark have been made publicly available, aiming to foster the development of more effective evaluation metrics within the field of medical informatics.
Our extensive experiments demonstrate the superiority of RaTEScore. We initially tested our metric against others on the public dataset ReXVal Yu et al. (2023a), and it shows superior alignment with human preference. Considering ReXVal’s limitation to chest X-ray reports, further testing was conducted on the diverse sub-tasks of RaTE-Eval, where RaTEScore consistently outperformed other metrics. We also conduct ablation studies to validate the effectiveness of different individual components in our pipeline.
@@ 2 Methods
In this section, we start by properly formulating the problem, and introducing the pipeline of our metric (Sec. 2.1). Then, we detail each of the modules for our metric computation, i.e., medical named entity recognition (Sec. 2.2), synonym disambiguation encoding (Sec. 2.3), and the final scoring procedure (Sec. 2.4). Lastly, we present the implementation details at each stage.
@@ 2.1 General Pipeline
Given two radiological reports, one is the ground truth for reference, denoting as $x$, and the other candidate for evaluation as $\hat{x}$. We aim to define a new similarity metric $S(x,\hat{x})$, that enables comparison of two radiological reports at the entity level, thus better reflecting their clinical consistency.
@@
$\mathbb{F}=\Phi_{\text{ENC}}(\Phi_{\text{NER}}(x))$ (1)
@@
$S(x,\hat{x})=\Phi_{\text{SIM}}(\mathbb{F},\mathbb{\hat{F}})$ (2)
@@ 2.2 Medical Named Entity Recognition
Overall, we categorize the entity types into five distinct groups within radiological contexts: {Anatomy, Abnormality, Disease, Non-Abnormality, Non-Disease}. Specifically, ‘Abnormality’ refers to notable radiological features such as masses, effusion, and edema. Conversely, ‘Non-Abnormality’ denotes cases where such abnormalities are negated in the context, as illustrated by the classification of ‘pleural effusion’ in the statement ‘No evidence of pleural effusion’. Compared to ‘Abnormality’, ‘Disease’ in radiology reports are more high-level, mainly about final main professional diagnosis conclusions, terms such as ‘pneumonia’ or ‘lymphadenopathy’.
@@
RaTE-NER Dataset. To support the development of our medical entity recognition module, we have constructed the RaTE-NER dataset, a large-scale, radiological named entity recognition (NER) dataset, including 13,235 manually annotated sentences from 1,816 reports within the MIMIC-IV database, that spans 9 imaging modalities and 23 anatomical regions, ensuring comprehensive coverage. Given that reports in MIMIC-IV are more likely to cover common diseases, and may not well represent rarer conditions, we further enriched the dataset with 33,605 sentences from the 17,432 reports available on Radiopaedia Rad , by leveraging GPT-4 and other medical knowledge libraries to capture intricacies and nuances of less common diseases and abnormalities. More details can be found in the Appendix A.2. We manually labeled 3,529 sentences to create a test set. As shown in Table 2 and Table 1, the RaTE-NER dataset offers a level of granularity not seen in previous datasets, with comprehensive entity annotations within sentences, that enables to train models for medical entity recognition within our analytical pipeline.
@@ 2.4 Scoring Procedure
$S(x,\hat{x})=\frac{\sum_{j}{W(t_{i^{*}},t_{j}})\cdot\text{sim}(e_{i^{*}},\hat{e}_{j})}{\sum_{j}{W(t_{i^{*}},t_{j})}}$ (4)
@@
$\text{sim}(e_{i},\hat{e}_{j})=\left\{\begin{aligned} {p}\cos(f_{i},\hat{f}_{j}),\quad if\quad t_{i}\neq t_{j}\\ \cos(f_{i},\hat{f}_{j}),\quad if\quad t_{i}=t_{j}\end{aligned}\right.$ (5)
@@
Finally, due to the order of performing max indexing for selecting referenced entities and weighted sum pooling on all candidate entities, the final similarity metric $S(x,\hat{x})$ does not comply with the commutative law. $S(x,\hat{x})$ and $S(\hat{x},x)$ can be analogous to precision and recall respectively. Thus, our final RaTEScore is defined as the harmonic mean of $S(x,\hat{x})$ and $S(\hat{x},x)$, following classical F1-score format:
$\texttt{RaTEScore}=\begin{cases}0,\quad\;\ \text{if }S(x,\hat{x})+S(\hat{x},x)=0,\\ 2\times\frac{S(x,\hat{x})\times S(\hat{x},x)}{S(x,\hat{x})+S(\hat{x},x)},\quad\ \text{otherwise}.\end{cases}$
@@ 3 RaTE-Eval Benchmark
Sentences-level Human Rating. Existing studies have predominantly utilized the ReXVal dataset Yu et al. (2023b), which requires radiologist annotators to identify and count errors in various potential categories. The metric’s quality is assessed by the Kendall correlation coefficient between the total number of errors and the result from the automatic metric. The possible error categories are as follows:
@@ 5. Mention the comparison that is absent from the reference impression;
Building on this framework, we introduce two improvements to enhance the robustness and applicability of our benchmark: (1) normalization of error counts: recognizing that a simple count of errors may not fairly reflect the informational content in sentences, we have adapted the scoring to annotate the number of potential errors. Specifically, we computed the sum of both correct and incorrect findings in the reference sentence. This sum represents the total number of potential errors that could occur. This approach normalizes the counts, ensuring a more balanced assessment across varying report complexities. (2) diversification of medical texts: unlike existing benchmarks that are limited to Chest x-rays from the MIMIC-CXR dataset Johnson et al. (2019), our dataset includes 2215 reports spanning 9 imaging modalities and 22 anatomies from the MIMIC-IV dataset Johnson et al. (2020), the involving imaging modalities and anatomies are listed in Appendix A.3.
@@
Paragraph-level Human Rating. Given that medical imaging interpretation commonly involves the evaluation of lengthy texts rather than isolated sentences, we also incorporate paragraph-level assessments into our analysis, from the MIMIC-IV reports. However, as it is challenging for humans to completely count all errors in long paragraphs accurately, we established a 5-point scoring system for our evaluations, following the RadPEER Goldberg-Stein et al. (2017) - an internationally recognized standard for radiologic peer review. They defines "Concur with Interpretation" as "correct diagnosis". In our annotation process, due to the difficulty of counting errors in paragraph-level reports (too detailed and some errors are not of great clincial significance), we instructed the radiologists to approximate the ratio of correct diagnosess based on their clinical judgement, following the ‘RadPEER’ standard, which is more aligned with the human rating system for report writing in clinical.
The scores range from 5, denoting a perfectly accurate report, to 0, that indicates the report without any correct observations. Detailed scoring criteria are provided in Appendix A.4, guiding radiologists on how to assign scores at different levels.
@@ 4.4.2 Entity Encoding Module Discussion
For the evaluation of Entity Encoding Module, we compare several off-the-shelf entity encoding models trained using different approaches on the sentence-level correlation task of RaTE-Eval. BioLORD-2023-C Remy et al. (2024) is trained on medical entity concepts, MedCPT-Query-Encoder Jin et al. (2023) is trained on PubMed user click search logs, while RadBERT Chambon et al. (2023), CXR-BERT Boecking et al. (2022), and BioViL-T Bannur et al. (2023) are pre-trained on a large corpus of radiology texts. As shown in Table 8, BioLORD, due to its original training goal covering medical entity normalization, which aligns with our needs in the Entity Encoding module, achieved the best performance. Based on these results, we selected BioLORD-2023-C as the base model for our Entity Encoding Module.
BioLORD MedCPT RadBERT CXR-BERT BioViL-T
Pearson $\tau$ 0.540 0.498 0.519 0.368 0.465
Table 8: Ablation Study on Pretrained Model of Entity Encoding Module.
@@ 5.2 Radiological Text Evaluation Metric
With the advancement of medical imaging analysis, researchers have recognized the importance of evaluating the quality of radiology text generation. Metrics such as CheXbert F1 Smit et al. (2020) and RadGraph F1  Yu et al. (2023a) are based on medical entity extraction models. However, CheXbert can only annotate 14 chest abnormalities, and RadGraph F1 Jain et al. is only trained on chest X-ray modality. MEDCON Yim et al. (2023) expands the extraction range by QuickUMLS package Soldaini and Goharian (2016), which relies on a string match algorithm that is not flexible. RadCliQ Yu et al. (2023a) performs ensembling with BLEU, BERTScore, CheXbert vector similarity, and RadGraph F1 for a comprehensive yet less interpretable evaluation. These metrics calculate the overlap between reference and candidate sentences while overlooking the issue of synonymy. Recently, metrics using Large Language Models (LLMs) such as GPT-4, such as G-Eval Liu et al. (2023), LLM-as-a-Judge Zheng et al. (2024), and LLM-RadJudge Wang et al. (2024) have emerged, closely mimic human evaluation levels. However, these methods are unexplainable and may have potential subjective bias. Besides, their high computational cost also limits them for statistic robust large-scale evaluation.
@@ 5.3 Medical Named-Entity Recognition
The MedNER task targets extracting medical-related entities from given contexts. Great efforts have been made in this domain Jin et al. (2023); Monajatipoor et al. (2024); Keloth et al. (2024); Li and Zhang (2023); Chen et al. (2023). Inspired by the success of this work, we believe MedNER models are strong enough to simplify and structure complex clinical texts, thus reducing the difficulty of automatically comparing two clinical texts. The most related work to ours is RadGraph Jain et al. which trained an NER model for Chest X-ray reports while we are targeting more general clinical reports regardless of their type.
@@ 6 Conclusion
Although our proposed metric, RaTEScore, has performed well across various datasets, there are still some limitations. First, in the synonym disambiguation module, we evaluated the performance of several existing models and directly ultilized them without fine-tuning specifically for the evaluation scenario, which could be enhanced in the future. Furthermore, while we expanded from single-modality radiological report evaluation to multimodal whole-body imaging, we still only considered the issues within the radiological report scenario and did not extend to other medical contexts beyond radiology, nor to the evaluation of other medical tasks, like medical QA, summarization task. These areas require more exploration.
Acknowledgements
This work is supported by the National Key R&D Program of China (No. 2022ZD0160702).
@@ A.2 Automatic Annotation Approach
Specifically, we manually annotate several reports at the required granularity and adopt few-shot prompts with GPT-4 to initially establish an NER dataset.
GPT-4 prompt: You are an AI assistant specializing in radiology reports reading. You are provided with a medical caption. Extract the entities and decide their type from organ, abnormal description or disease. Collect the organ and description together if the description modifies the organ. Leave disease alone. Make sure that the description is about the abnormality but not position.
The output should follow this format: [organs; abnormal description] or [disease]. All words in [] should belong to the original sentence.
Few-shot examples: ’context’: "The sentence is: Hetergeneous and nodular enhancement of the liver with pre-contrast HU of -4 (!) indicating hepatic steatosis.", ’response’: " [ liver; Hetergeneous and nodular enhancement ] [ liver; pre-contrast HU of -4 ] [ hepatic steatosis ] "
Following this, we build a robust medical entity library, integrating UMLS Bodenreider (2004), Snomed CT Donnelly et al. (2006), ICD-10 ICD , and other knowledge bases, then, compare all extracted entities using the MedCPT Jin et al. (2023) model for similarity. During the comparison process, entities with cosine similarity lower than 0.83 were filtered out. Most entities below this threshold did not meet our requirements. Subsequently, we removed sentences with an entity annotation density lower than 0.7 at the sentence level. Finally, we use medspaCy Eyre et al. (2021) and also key negative words detection in reports, such as “no”, “without”, “unremarkable”, “intact”, to determine the positive or negative polarity of each word in the sentence.